After-Tax Cash Flows

Below is a list of aspects of various capital expenditure proposals that the capital budgeting team of Modern Systems, Inc., has incorporated into its net present value analyses during the past year. Unless otherwise noted, the items listed are unrelated to each other. All situations assume a 30% income tax rate and a 10% minimum desired rate of return.

1. Pre-tax savings of $5,000 in cash expenses will occur in each of the next three years.

2. A machine is purchased now for $82,000.

3. Special tools costing $54,000 will be depreciated $10,800, $21,600, and $21,600, respectively, on the tax return over a three-year life.

4. A patent purchased for $357,000 will be amortized on a straight-line basis over 15 years on the tax return. No salvage value is expected.

5. Pre-tax savings of $8,000 in cash expenses will occur in each of the next seven years.

6. Pre-tax savings of $8,500 in cash expenses will occur in the first, fourth, and seventh years from now.

7. The special tools described in aspect 3 will be sold after three years for $13,000 cash.

8. A truck with a tax book value of $7,600 after two years will be sold at that time for $4,600.

a. Calculate and record in column A the related after-tax cash flow effect(s).

b. Indicate in column B the timing of each cash flow shown in column A. Use 0 to indicate immediately and 1, 2, 3, 4, and so on for each year involved.

The answer to investment aspect 1 is presented as an example.

Use negative signs with answers that are cash outflows. Under Column B, select the appropriate year for the timing of each cash flow using the drop-down menu.
Answer #1

Dear student,

1. Post tax inflow = $5000(0.7)=$3500 per annum.

PV= $3500(2.487)= $ 8703. Where 2.487 is annuity for 3 yrs @10% return.

2. ($82000)*1= ($82000) i.e outflow in the zero th year

3. Special tool purchased ($54000) * 1= ($54000)

Tax savings [ ie multiply by 30%] on depreciation

$10800(0.3) * 0.9091 [ie PV factor of year 1] = 2945

$21600(0.3) * 0.8264 [ie PV factor of yr 2]= 5355

$21600(0.3) * 0.7513 [ie PV factor of yr 3] = 4868

Net= ($40832)

4. Patent purchased= $357000* 1[PV factor of yr 0] = ($357000)

Tax benefit on amortization over 15 yrs

$23800*0.3*7.6061= $54308. Where 0.3 is for tax amt. and 7.6061 represents annuity factor for 15 yrs @10% return

Net= ($302692)

5. Post tax savings = $8000(0.7)= $5600 per year

Hence PV of savings over 7 yrs = $5600*4.868 =$ 26241.6

6. Post tax = $8500*0.7= $5950

Hence yr 1=$5950*0.9091= $5409

Yr 4= $5950*0.6830= $4064

Yr 7= $5950*0.5132= $ 3054

Net= $12527

7. The special tool has been entirely depreciated hence entire salvage amt received shall be taxable as capital gain.

Post tax inflow= $13000* 0.7* 0.7513= $6834

8. Capital loss on sale =7600-4600= $3000

Tax benefit on capital loss= 3000*0.3= $ 900

Cash Inflow from sale =4600

Total inflow= $ 5500

Hope this helps...thank you!!